OpenStudy (stuck_in_the_middle):

A weight suspended from a length of string forms a pendulum. The amplitude, R is 0.35m and the periodic time is 4.5 seconds. Find the length of the piece of string and the maximum velocity and maximum acceleration. When the displacement x = 0.12m, what are the velocity and acceleration.

OpenStudy (anonymous):

period = 2 pi sqrt(L/g) from L and period get velocity as distance/time model : y = A sin (w t) = A sin(2 pi t/T) vel = dy/dt accel = (d/dt)(dv/dt)